Monitoring/Measuring T_{max}

1. Second-Derivative method.

• Find the ‘cusp’ in the X − T correlation.

• Fails miserably; uncertainties are too large.

2. Set a threshold on the X − T correlation.

• Find the drift time when the X − T correlation reaches X_{max}.

• Works reasonably well.

3. Set a threshold on the integrated drift time spectrum.

• The integrated drift time spectrum is proportional to X(T).

• Find the fraction of the integrated sum that returns an accurate T_{max}

• Works better than 2 (smaller uncertainty).

Results of using a threshold set on the integrated drift time spectrum to measure T_{max} (run 8774).

Region Method 2 T_{max} (ns) Method 3 T_{max} (ns) Qin, et al T_{max} (ns)
1  198 ± 11  216 ± 1  195
2  591 ± 76  463 ± 2  526
3  1144 ± 71  1144 ± 4  1150

Comparison of T_{max} measured with different methods.

Results of measurements of T_{max} for several runs using a threshold set on the integrated drift time spectrum.

Monitoring/Measuring T_{0}

1. Fit the small X drift time spectrum.

• Use functional form from study of discrete ionization.

• Fit parameters are ambiguous.

2. Set a threshold on the small X drift time spectrum. • Find reasonable lower limit.

• It works.

Results of measuring T_{0} using a threshold applied to the small-X drift time spectrum for each region.

Results of measurements of T_{0} for several runs using a threshold set on the small-X drift time spectrum.